# Aggregation Schemes for Single-Vector WSI Representation Learning in Digital Pathology

### v1

@@ Abstract
Histopathology, Whole Slide Image, WSI, Image Representation, WSI Search, Set Representation Learning, Single-Vector, Aggregation.
I Introduction
@@
The image analysis field has witnessed remarkable advancements in recent years, thanks to the integration of deep learning models and the availability of large-scale data. However, analysis of histopathology images using deep learning is not as straightforward as ordinary-sized images. Especially, to build whole slide image (WSI) search engines, the most critical step is to be able to measure the similarity between WSIs efficiently and accurately [Tizhoosh and Pantanowitz, 2024, Lahr et al., 2024]. The state-of-the-art approach to this step is to extract high-quality deep embeddings that capture tissue morphology from each WSI using a well-trained deep model. To this end, it is necessary to feed WSIs to build such models. However, given the gigapixel nature of WSIs, feeding a WSI at high magnification into a model to GPU memory is infeasible. To overcome this challenge, a WSI is usually split into a set of much smaller sub-images called “patches” or tiles. Following this procedure, we end up with a set of deep embeddings per WSI which makes it challenging to employ them for downstream tasks like WSI retrieval (a set of pacthes that could be very large if a smart patch selection is not available). In this situation, having one single-vector WSI embedding can mitigate such challenges. With single-vector WSI representation much less memory is required to build a WSI search system. As well, using distance measures calculating similarity between WSIs becomes trivial. Finally, the search speed will increased and storage requirements will decrease resulting in wider deployment of WSI image classification and retrieval.
@@
A practical and established approach to calculate the similarity between two sets of embeddings is the median-of-minimums, a scheme that was introduced as part of the Yottixel search engine [Kalra et al., 2020b]; one calculates the minimum distance of any given input patch to all other patches (of the other WSI) and takes the median of all minimum distances overall (see Fig. 1). This approach is quite practical and fast (as Yottixel used binary feature vectors) but it is not able to provide a single vector to represent WSI. A more desirable solution would be to obtain one single embedding per WSI for more efficient storage and more targeted WSI-to-WSI matching, or for WSI classification (Fig. 2).
@@
To this end, different aggregation algorithms to derive a single vector of deep features from a set of patch embeddings have been proposed including simple average or max pooling operations, Deep Sets [Zaheer et al., 2017], memory networks [Kalra et al., 2020a], focal attention [Kalra et al., 2021], graph convolution neural networks [Adnan et al., 2020], Gaussian mixture model (GMM) Fisher Vector [Jaakkola and Haussler, 1999, Perronnin and Dance, 2007, Hemati et al., 2023], and Deep Fisher Vector and its variations (binary and sparse Fisher Vector) [Hemati et al., 2023].
@@
Figure 1: Yottixel [Kalra et al., 2020b], although a patch-oriented search engine, introduced the first practical approach for WSI-to-WSI comparison through the median-of-minimum distance measurements.
@@
II Background
@@ II-A GMM Fisher Vector (1999-2007)
Fisher Vector can be seen as an extension to bag of visual words (BoVW) [Csurka et al., 2004] that was initially developed in computer vision research for encoding a set of local image descriptors into one embedding. Theoretically, Fisher Vector can be developed on top of any generative models [Jaakkola and Haussler, 1999]. Some works [Perronnin and Dance, 2007] have introduced Gaussian mixture model (GMM)-based Fisher Vector which can be calculated using the gradient of the log-likelihood of the GMMs with respect to its parameters given a bag of data points (e.g., means, variances, and mixing coefficients). Theoretically, while BoVW only employs mean statistics to obtain the set representation, in the GMM-based Fisher Vector, the variance as a higher-order statistic is also used.
II-B Deep Sets (2017-2021)
@@
The Deep Set approach [Zaheer et al., 2017] focuses on permutation invariant representation learning and establish a standard definition of the permutation invariant functions. This approach shows such a family of functions can be implemented as neural networks to learn permutation invariant representations. Other works employ Deep Sets-like architecture along with CNN to learn end-to-end WSI representations in WSI search tasks [Hemati et al., 2021]. These methods showed the obtained representations achieve better search performance both in terms of search speed and accuracy compared with Yottixel search engine [Kalra et al., 2020b].
II-C Memory Networks (2020)
Other works have presented Memory-based Exchangeable Model (MEM) to learn the set functions [Kalra et al., 2020a]. MEM employs memory units and self-attention mechanisms to capture inter-dependencies between instances by mapping the input sets to high-level features. MEM was evaluated on point cloud and lung WSIs classification.
II-D Focal Attention (2021)
Another set representation scheme developed for WSI classification/search is Focal attention [Kalra et al., 2021]. This approach was inspired by two recent developments in representation learning literature, namely focal loss and attention mechanism. More precisely, the focal attention approach maps embeddings for all patches of a mosaic into a single embedding using an attention-weighted averaging modulated by a trainable focal factor.
II-E Deep Fisher Vector and its variations (2023)
@@
Motivated by the success of GMM-based Fisher Vector, Fisher Vector theory has been proposed for more recent deep generative models including Variational Autoencoders (VAEs) [Qiu et al., 2017] and GANs [Zhai et al., 2019]. Recently, other works have extended VAE-based Fisher Vectors for the WSI representation learning task [Hemati et al., 2023]. More specifically, this approach offered two variations of Deep Fisher Vectors, namely “deep sparse Fisher Vector” and “deep binary Fisher Vector”, to obtain binary and sparse permutation-invariant WSI embeddings suitable for downstream tasks such as efficient WSI search. To obtain the deep sparse and binary Fisher Vectors, one can employ a simple VAE where its latent space is also connected to a dense layer with a softmax activation function that outputs a probability distribution representing the predicted WSI class. This classifier injects class information (i.e., diagnosis) into the final WSI embeddings. Considering that the Fisher Vector is derived from the gradient space of the VAE, to achieve sparse or binary embeddings, it is desired to have sparse or minimum quantization loss gradients. This step has been performed with a gradient regularization loss which is calculated using double backpropagation [Hemati et al., 2023]. Finally, the total loss for training the VAE consists of a weighted summation of reconstruction loss, KL divergence loss, classification loss, and gradient regularization loss. The contribution of gradient regularization loss to the total loss is controlled through the $\alpha$ regularization parameter. Fig. 3 shows the training architecture.
@@
After training the VAE, first, the gradients of the reconstruction loss are computed. Second, the gradient using power and $L_{2}$ normalization steps are normalized (see [Perronnin et al., 2010]). Finally, considering that the dimensionality of the deep Fisher Vector is the same as the number of the VAE parameters (which can be a large number making the WSI search expensive), the effective dimensions are selected to be the top $M$ parameters that provide the highest variance in their respective gradient values for the training data. Fig. 4 shows the calculation of single-vector WSI representation given the trained VAE and patch embeddings for a WSI.
@@
III Methodology
@@
III-A Datasets
The datasets presented in this study contain a collection of four histopathology cases, including bladder, breast, kidney, and colon primary sites disease cases. All datasets are drawn from The Cancer Genome Atlas (TCGA). For each WSI, multiple patches were extracted using Yottixel’s mosaic algorithm (a two-staged clustering approach that employs stain histograms and patch proximity to group patches) [Kalra et al., 2020b]. Each patch is represented by an embedding extracted from DenseNet [Huang et al., 2017] with imageNet pre-trained weights encoding the structural information and unique characteristics of the corresponding histological region. Patch embeddings enable quantitative analysis, facilitating comparisons among different WSI regions providing a valuable resource for exploring morphology matching, developing image analysis algorithms, and advancing diagnostic and prognostic research in oncology.
@@ III-A Datasets
TABLE I: Bladder Metrics
@@
TABLE II: Breast Metrics
@@
TABLE III: Kidneys Metrics
@@
TABLE IV: Colon Metrics
III-B Benchmarking Scheme and Baselines
WSI search and retrieval offers many benefits including teleconsultation, reduced workload, improved diagnostic quality, and expedited adoption and democratization of digital pathology through more efficient indexing of tissue images [Tizhoosh and Pantanowitz, 2018, Janowczyk and Madabhushi, 2016, Lahr et al., 2024, Tizhoosh and Pantanowitz, 2024]. Considering these benefits, we are motivated to investigate the quality of WSI embeddings through the lens of $k$-Nearest Neighbors ($k$-NN) WSI search and retrieval.
The accuracy, Macro F1 score, and weighted F1 score performance measures are used as the evaluation metrics. We benchmark multiple techniques against Yottixel’s “median of minimums” as a non-aggregation approach [Kalra et al., 2020b]. We test max and mean poolings, Gaussian Mixture Model-based Fisher Vector [Sánchez et al., 2013], Deep Sets [Zaheer et al., 2017], memory networks [Kalra et al., 2020a], focal attention [Kalra et al., 2021], and deep sparse Fisher Vector and deep binary Fisher Vector by [Hemati et al., 2023].
@@ III-B Benchmarking Scheme and Baselines
IV Results
Tables I, II, III and IV show the evaluation results for bladder, breast, kidney, and colon datasets where DenseNet was used as feature extractors. Clearly, in most cases, Deep Fisher Vector and its variations (sparse and binary) achieve the best overall performance compared with all other baselines. Here we should emphasize sparse embeddings can reduce the memory required to store embeddings while binary WSI embeddings provide both memory efficiency and extremely fast search speed by employing Hamming distance. As a sample case, we also measured the time required to conduct the colon dataset search experiment for both Deep Sparse and Binary Fisher Vectors and also the median of minimums. For the Sparse Fisher Vector, Euclidean distance is used while for the Binary Fisher Vector Hamming distance is employed. The results for this experiment are presented in Table V. Clearly, the Deep Binary Fisher Vector achieves the fastest search speed. Given that for the Sparse Fisher Vector the Euclidean distance is used, its search speed is the same as other methods that output one real-valued embedding (with the same dimension) per WSI.
@@ IV Results
TABLE V: Time required for different dimensions of Fisher vector for Sparse and Binary methods for the Colon dataset compared to Yottixel
IV-A Ablation study
To study the effect of the gradient sparsity and quantization losses for Deep Sparse and Binary Fisher vector embeddings respectively, we performed an ablation study on $\alpha$ for skin and lung datasets. To this end, we train the VAE for different values of $\alpha$, while keeping all other hyperparameters the same. The results for breast and colon datasets are presented in Tables VI and VII. Clearly, including the regularization loss further improves the quality of embeddings while increasing memory and speed efficiency by introducing binary nature and sparsity characteristics for WSI embedding.
@@ IV-A Ablation study
TABLE VI: Results for Sparse and Binary Fisher vector methods for Breast dataset with different values of $\alpha$, including Accuracy, Macro $\bar{F}1$, and Weighted $\bar{F}1$
@@
TABLE VII: Results for Sparse and Binary Fisher vector methods for Colon dataset with different values of $\alpha$, including Accuracy, Macro $\bar{F}1$, and Weighted $\bar{F}1$
V Conclusion

### v2

# Mayo Clinic, Rochester, MN, USA

@@ Abstract
Histopathology Whole Slide Image WSI Image Representation WSI Search Set Representation Learning Single-Vector Aggregation
1 Introduction
@@
The image analysis field has witnessed remarkable advancements in recent years, thanks to the integration of deep learning models and the availability of large-scale data. However, analysis of histopathology images using deep learning is not as straightforward as ordinary-sized images. Especially, to build whole slide image (WSI) search engines, the most critical step is to be able to measure the similarity between WSIs efficiently and accurately [1, 2]. The state-of-the-art approach to this step is to extract high-quality deep embeddings that capture tissue morphology from each WSI using a well-trained deep model. To this end, it is necessary to feed WSIs to build such models. However, given the gigapixel nature of WSIs, feeding a WSI at high magnification into a model to GPU memory is infeasible. To overcome this challenge, a WSI is usually split into a set of much smaller sub-images called “patches” or tiles. Following this procedure, we end up with a set of deep embeddings per WSI which makes it challenging to employ them for downstream tasks like WSI retrieval (a set of pacthes that could be very large if a smart patch selection is not available). In this situation, having one single-vector WSI embedding can mitigate such challenges. With single-vector WSI representation much less memory is required to build a WSI search system. As well, using distance measures calculating similarity between WSIs becomes trivial. Finally, the search speed will increased and storage requirements will decrease resulting in wider deployment of WSI image classification and retrieval.
@@
A practical and established approach to calculate the similarity between two sets of embeddings is the median-of-minimums, a scheme that was introduced as part of the Yottixel search engine [3]; one calculates the minimum distance of any given input patch to all other patches (of the other WSI) and takes the median of all minimum distances overall (see Fig. 1). This approach is quite practical and fast (as Yottixel used binary feature vectors) but it is not able to provide a single vector to represent WSI. A more desirable solution would be to obtain one single embedding per WSI for more efficient storage and more targeted WSI-to-WSI matching, or for WSI classification (Fig. 2).
@@
To this end, different aggregation algorithms to derive a single vector of deep features from a set of patch embeddings have been proposed including simple average or max pooling operations, Deep Sets [4], memory networks [5], focal attention [6], graph convolution neural networks [7], Gaussian mixture model (GMM) Fisher Vector [8, 9, 10], and Deep Fisher Vector and its variations (binary and sparse Fisher Vector) [10].
@@
Figure 1: Yottixel [3], although a patch-oriented search engine, introduced the first practical approach for WSI-to-WSI comparison through the median-of-minimum distance measurements.
@@
1.1 Statement of Significance
Whole Slide Image (WSI) representation learning is a cornerstone of computational pathology, enabling efficient retrieval, classification, and analysis of gigapixel-scale histopathology images. However, the challenge lies in deriving a single high-quality feature vector for an entire WSI, given that existing GPUs cannot process them as single images due to their enormous resolution. Traditionally, WSIs are divided into smaller patches, each independently embedded using deep neural networks, leading to a set-based representation problem where optimal aggregation strategies remain an open question. This study systematically evaluates multiple state-of-the-art set representation learning techniques—including Deep Sets, Memory Networks, Focal Attention, Gaussian Mixture Model (GMM) Fisher Vector, and deep sparse and binary Fisher Vector—against conventional pooling methods for WSI embedding aggregation. Additionally, we benchmark these approaches against a non-aggregating retrieval technique based on the median of minimum distances of patch embeddings. Our comprehensive evaluation, conducted on TCGA datasets across four cancer types (bladder, breast, kidney, and colon), provides critical insights into the effectiveness of different aggregation strategies for WSI search and retrieval. By rigorously comparing various WSI embedding aggregation techniques, this study advances the field of computational pathology by guiding researchers and practitioners toward more effective and computationally feasible WSI retrieval methodologies. The findings have significant implications for large-scale histopathology archives, image-based precision medicine, and AI-driven diagnostic workflows.
2 Background
@@ 2.1 GMM Fisher Vector (1999-2007)
Fisher Vector can be seen as an extension to bag of visual words (BoVW) [11] that was initially developed in computer vision research for encoding a set of local image descriptors into one embedding. Theoretically, Fisher Vector can be developed on top of any generative models [8]. Some works [9] have introduced Gaussian mixture model (GMM)-based Fisher Vector which can be calculated using the gradient of the log-likelihood of the GMMs with respect to its parameters given a bag of data points (e.g., means, variances, and mixing coefficients). Theoretically, while BoVW only employs mean statistics to obtain the set representation, in the GMM-based Fisher Vector, the variance as a higher-order statistic is also used.
2.2 Deep Sets (2017-2021)
@@
The Deep Set approach [4] focuses on permutation invariant representation learning and establishes a standard definition of the permutation invariant functions. This approach shows such a family of functions can be implemented as neural networks to learn permutation invariant representations. Other works employ Deep Sets-like architecture along with CNN to learn end-to-end WSI representations in WSI search tasks [12]. These methods showed the obtained representations achieve better search performance both in terms of search speed and accuracy compared with Yottixel search engine [3].
2.3 Memory Networks (2020)
Other works have presented Memory-based Exchangeable Model (MEM) to learn the permutation invariant set functions [5]. In this method, a MEM is constructed from multiple memory blocks where each memory block is a sequence-to-sequence model containing multiple memory units. The output of each memory block is invariant to the permutations of the input sequence. The memory units are in charge of calculation of attention values for each instance. To calculate these attention values, the memory vectors are aggregated using a pooling operation (weighted averaging) to form a permutation-invariant representation. Increasing the number of memory units enables the memory block to capture more complex dependencies between the instances of a set by providing an explicit memory representation for each instance in the sequence. MEM was evaluated on point cloud and lung WSIs classification.
2.4 Focal Attention (2021)
Another set representation scheme developed for WSI classification/search is Focal attention [6]. This approach was inspired by two recent developments in representation learning literature, focal loss and attention based MIL [13] and proposed the novel pooling layer called focal attention (FocAtt-MIL) where all patches of a mosaic are mapped into a single embedding using an attention-weighted averaging [13] modulated by a trainable focal factor. More precisely, FocAtt-MIL is composed from four main components. Prediction MLP, WSI Context, Attention Module, and Focal Network. The Prediction MLP is a trainable neural network that calculates a prediction for each patch embedding in the set. WSI context is also a neural network (basically Deep Sets from [4] that obtains one embedding per WSI in a simple efficient manner that capture general information from the WSI. The attention module is made from two MLPs transformation, and the Attention networks. The attention module takes the patch embedding and WSI context and output the an attention value between 0 to 1 for each instance. Finally, the focal network receive the WSI context and computes a focal factor per dimension which further guides the final prediction toward better WSI representation.
2.5 Deep Fisher Vector and its variations (2023)
@@
Motivated by the success of GMM-based Fisher Vector, Fisher Vector theory has been proposed for more recent deep generative models including Variational Autoencoders (VAEs) [14] and GANs [15]. Recently, other works have extended VAE-based Fisher Vectors for the WSI representation learning task [10]. More specifically, this approach offered two variations of Deep Fisher Vectors, namely “deep sparse Fisher Vector” and “deep binary Fisher Vector”, to obtain binary and sparse permutation-invariant WSI embeddings suitable for downstream tasks such as efficient WSI search. To obtain the deep sparse, and binary Fisher Vectors, one can employ a simple VAE where its latent space is also connected to a dense layer with a softmax activation function that outputs a probability distribution representing the predicted WSI class. This classifier injects class information (i.e., diagnosis) into the final WSI embeddings. Considering that the Fisher Vector is derived from the gradient space of the VAE, to achieve sparse or binary embeddings, it is desired to have sparse or minimum quantization loss gradients. This step has been performed with a gradient regularization loss which is calculated using double backpropagation [10]. Finally, the total loss for training the VAE consists of a weighted summation of reconstruction loss, KL divergence loss, classification loss, and gradient regularization loss. The contribution of gradient regularization loss to the total loss is controlled through the $\alpha$ regularization parameter. Fig. 3 shows the training architecture.
@@
After training the VAE, first, the gradients of the reconstruction loss are computed. Second, the gradient using power and $L_{2}$ normalization steps are normalized (see [16]). Finally, considering that the dimensionality of the deep Fisher Vector is the same as the number of the VAE parameters (which can be a large number making the WSI search expensive), the effective dimensions are selected to be the top $M$ parameters that provide the highest variance in their respective gradient values for the training data. Fig. 4 shows the calculation of single-vector WSI representation given the trained VAE and patch embeddings for a WSI.
@@
2.6 Other Approaches
Recently foundation models for multiple WSI representation learning have been proposed that use various techniques for aggregation of patch embedding, including GIGAPATH [17], PRISM [18], CHIEF [19], and THREADS [20]. Among these works, GIGAPATH [17] and THREADS [20] require embeddings for all patches from WSI to perform the aggregation. Given that the backbones are generally foundation models, such techniques are extremely expensive in practice. Additionally, CHIEF [19] and PRISM [18] use techniques based on the attention-based MIL by [13]. Considering that here we already discussed two different extensions of the attention-based MIL namely Focal attention [6] and Memory-based Exchangeable Model [5] techniques, we did not include the aggregation techniques in CHIEF [19] and PRISM.
In terms of acquiring expressive embeddings, foundation models have been reported to be inaccurate for zero-shot WSI retrieval [21], insensitive to inter-center variations [22], and exhibit low rotational invariance [23]. Moreover, they consume approximately 35 times more energy than conventional deep models, are not easily fine-tunable, and still require large amounts of data for downstream tasks [24].
3 Methodology
@@
3.1 Datasets
The datasets presented in this study contain a collection of four histopathology cases, including bladder, breast, kidney, and colon primary sites disease cases. All datasets are drawn from The Cancer Genome Atlas (TCGA). For each WSI, multiple patches were extracted using Yottixel’s mosaic algorithm (a two-staged clustering approach that employs stain histograms and patch proximity to group patches) [3]. Each patch is represented by an embedding extracted from DenseNet [25] with imageNet pre-trained weights encoding the structural information and unique characteristics of the corresponding histological region.
@@ 3.1 Datasets
Table 1: Bladder Metrics
@@
Table 2: Breast Metrics
@@
Table 3: Kidneys Metrics
@@
Table 4: Colon Metrics
3.2 Benchmarking Scheme and Baselines
WSI search and retrieval offers many benefits including teleconsultation, reduced workload, improved diagnostic quality, and expedited adoption and democratization of digital pathology through more efficient indexing of tissue images [26, 27]. Considering these benefits, we are motivated to investigate the quality of WSI embeddings through the lens of $k$-Nearest Neighbors ($k$-NN) WSI search and retrieval.
The accuracy, Macro F1 score, and weighted F1 score performance measures are used as the evaluation metrics. We benchmark multiple techniques against Yottixel’s “median of minimums” as a non-aggregation approach [3]. We test max and mean poolings, Gaussian Mixture Model-based Fisher Vector [28], Deep Sets [4], memory networks [5], focal attention [6], and deep sparse Fisher Vector and deep binary Fisher Vector by [10].
@@ 3.2 Benchmarking Scheme and Baselines
4 Results
Tables 1, 2, 3 and 4 show the evaluation results for bladder, breast, kidney, and colon datasets where DenseNet was used as feature extractors. Clearly, in most cases, Deep Fisher Vector and its variations (sparse and binary) achieve the best overall performance compared with all other baselines. Here we should emphasize sparse embeddings can reduce the memory required to store embeddings while binary WSI embeddings provide both memory efficiency and extremely fast search speed by employing Hamming distance. As a sample case, we also measured the time required to conduct the colon dataset search experiment for both Deep Sparse and Binary Fisher Vectors and also the median of minimums. For the Sparse Fisher Vector, Euclidean distance is used while for the Binary Fisher Vector Hamming distance is employed. The results for this experiment are presented in Table 5. Clearly, the Deep Binary Fisher Vector achieves the fastest search speed. Given that for the Sparse Fisher Vector the Euclidean distance is used, its search speed is the same as other methods that output one real-valued embedding (with the same dimension) per WSI.
@@ 4 Results
Table 5: Time required for different dimensions of Fisher vector for Sparse and Binary methods for the Colon dataset compared to Yottixel
4.1 Ablation study
To study the effect of the gradient sparsity and quantization losses for Deep Sparse and Binary Fisher vector embeddings respectively, we performed an ablation study on $\alpha$ for skin and lung datasets. To this end, we train the VAE for different values of $\alpha$, while keeping all other hyperparameters the same. The results for breast and colon datasets are presented in Tables 6 and 7. Clearly, including the regularization loss further improves the quality of embeddings while increasing memory and speed efficiency by introducing binary nature and sparsity characteristics for WSI embedding.
@@ 4.1 Ablation study
Table 6: Results for Sparse and Binary Fisher vector methods for Breast dataset with different values of $\alpha$, including Accuracy, Macro $\bar{F}1$, and Weighted $\bar{F}1$
@@
Table 7: Results for Sparse and Binary Fisher vector methods for Colon dataset with different values of $\alpha$, including Accuracy, Macro $\bar{F}1$, and Weighted $\bar{F}1$
5 Conclusion